Calculate the number of H+ ion present in 1 ml of a solution whose pH is 10.

Grade:

2 Answers

Sunil Kumar FP
9 years ago
For expressing hydrogen ion concentration we are taking th help of ph which is exxpressed as
ph=-log[H+]
let the number of H+ ion be x.since ph is the no of H+ion in 1 lt of the solution therefore it is given by the formula
10=-log[x/10^-3]
x=10^-13
no of H+ ions=10^-13*6.023*10^23
=6.023*10^10
Pratishtha Sahu
6 years ago
pH = 13[H+] = 10^-13 mol/lit since vol=1 ml = 10^-3 lit[H+]=10^-13 × 10^-3 = 10^-16 molsince 1 mol contain 6.022 × 10^23 H+ions therefore no of H+ ions = 6.022×10^-23×10^-16= 6.022×10^7